### v3

# Camouflaged Object Tracking: A Benchmark

Abstract
Visual tracking has seen remarkable advancements, largely driven by the availability of large-scale training datasets that have enabled the development of highly accurate and robust algorithms. While significant progress has been made in tracking general objects, research on more challenging scenarios, such as tracking camouflaged objects, remains limited. Camouflaged objects, which blend seamlessly with their surroundings or other objects, present unique challenges for detection and tracking in complex environments. This challenge is particularly critical in applications such as military, security, agriculture, and marine monitoring, where precise tracking of camouflaged objects is essential. To address this gap, we introduce the Camouflaged Object Tracking Dataset (COTD), a specialized benchmark designed specifically for evaluating camouflaged object tracking methods. The COTD dataset comprises 200 sequences and approximately 80,000 frames, each annotated with detailed bounding boxes. Our evaluation of 20 existing tracking algorithms reveals significant deficiencies in their performance with camouflaged objects. To address these issues, we propose a novel tracking framework, HiPTrack-MLS, which demonstrates promising results in improving tracking performance for camouflaged objects. COTD and code are avialable at https://github.com/openat25/HIPTrack-MLS.
Index Terms:
Visual tracking; Camouflaged object tracking; Benchmark
I Introduction
In recent years, the field of visual tracking has made significant advances, largely due to the widespread use of large-scale training datasets. The richness of these datasets has facilitated the development of high-precision and robust tracking algorithms [1, 2, 3, 4, 5, 6, 7, 8]. Although substantial progress has been made in general object tracking, research into more specialized and challenging scenarios remains limited. Among these, tracking camouflaged objects is particularly difficult because these targets seamlessly blend into their surroundings, making them hard to detect with the naked eye and significantly increasing tracking complexity [9, 10]. While existing tracking methods have achieved impressive results on regular objects through various techniques such as high-level semantics [11], block-based correlation filters [12], spatio-temporal feature learning [13], and long-term tracking frameworks [14], their performance degrades significantly when dealing with camouflaged targets due to the inherent visual ambiguity between the target and background.
(a) Example of generic object tracking.
(b) Example of camouflaged object tracking.
Figure 1: Generic object tracking (a) and camouflaged object tracking (b). Compared to generic object tracking, tracking camouflaged objects is more challenging due to their tendency to blend into the background, resulting in indistinguishable visual features.
Camouflaged objects play a crucial role in several critical applications across various domains. For instance, in the military and defense sector, the ability to identify and track camouflaged objects, such as hidden weapons or concealed enemy units, is vital for battlefield awareness and operational success [15, 16]. In the field of autonomous driving, camouflaged objects, such as animals or debris that blend into the road environment, pose significant challenges for vehicle perception systems [17, 18, 19]. In medical imaging, camouflaged objects often take the form of pathological regions, such as tumors that are difficult to distinguish from surrounding healthy tissue [20, 21]. Additionally, in wildlife monitoring and ecological research, camouflaged animals often need to be tracked without disturbing their natural behavior [22, 23, 24].
Distinguishing camouflaged objects from their backgrounds in these applications is crucial, but it involves significant challenges like background clutter, boundary blurring, texture mimicry, color similarity, varying lighting conditions, and occlusion [25, 9, 10]. Addressing these challenges through innovative approaches and advanced technologies can lead to significant advancements in applications involving camouflaged objects [26, 27, 28, 29, 30]. In recent years, numerous computer vision techniques have been developed to address the challenges posed by camouflaged objects [31, 30]. One prominent direction is Camouflaged Object Detection (COD), which identifies objects blending into their surroundings due to similar textures, shapes, or colors. COD locates these hidden objects within images or videos [32, 33]. Camouflaged Object Segmentation (COS) takes this further by not only detecting but also accurately separating these objects from their backgrounds. COS involves classifying each pixel to determine if it belongs to the object or the background, offering detailed information about the object’s shape and edges [34]. The ongoing research in COD and COS continue to push the boundaries, promising further innovations and applications that tackle real-world challenges related to camouflaged objects.
Advancements in COD and COS technologies have been significantly propelled by the integration of deep learning algorithms. Methods such as MirrorNet [34], PFNet [25], and S-MGL [35] have greatly improved detection and segmentation accuracy by leveraging vast datasets to learn and generalize the subtle differences between camouflaged objects and their backgrounds. To support the development and evaluation of these algorithms, large-scale benchmark datasets like NC4K [30] and COD10K [32] have been established, providing crucial resources for advancing the field. Despite these advancements, Camouflaged Object Tracking (COT), which focuses on maintaining consistent identification of camouflaged objects throughout a video while accounting for changes in position, scale, and appearance, remains under-explored. This task requires not only the detection of camouflaged targets but also the continuous tracking of their location and status across video streams. To achieve this, more refined visual feature extraction and sophisticated tracking algorithms are necessary to prevent target loss or misalignment. Additionally, the inherent uncertainty and instability of target motion and the presence of coexisting objects further complicate the tracking process [36, 37, 38]. These challenges highlight the need for substantial innovation and effort to effectively address the complexities of COT in real-world scenarios. However, there is a critical gap in the field: to our knowledge, no specialized dataset exists for camouflaged object tracking. This absence significantly limits the development and rigorous evaluation of COT algorithms, as existing datasets may not adequately represent the unique challenges posed by camouflaged objects. Without a dedicated dataset, researchers face difficulties in testing their models against the specific nuances of camouflaged objects.
Figure 2: Some examples of annotations for COTD. Each sequence is annotated with axis-aligned bounding boxes and attributes.
In this work, we take the first step in addressing the challenging task of tracking camouflaged objects by introducing the Camouflaged Object Tracking Dataset (COTD). COTD features a diverse range of camouflaged objects across various environments, specifically designed to push the boundaries of camouflaged object tracking algorithms. The dataset includes 200 sequences totaling around 80,000 frames, with each sequence meticulously annotated with detailed bounding boxes and relevant attributes for comprehensive performance evaluation and analysis. To the best of our knowledge, COTD is the first benchmark dataset specifically dedicated to the task of tracking camouflaged objects. Fig. 1 demonstrates the differences between generic object tracking and camouflaged object tracking. Compared with generic object tracking, where targets are more distinguishable from their backgrounds, tracking camouflaged objects is notably more challenging as these objects blend seamlessly with their surroundings. This blending creates ambiguous visual cues that significantly complicate the tracking process. Our extensive dataset captures a broad spectrum of real-world scenarios involving camouflaged objects, laying a solid foundation for developing and evaluating new tracking algorithms that adapt to these conditions. We also conducted extensive evaluations of 20 state-of-the-art tracking algorithms and found that current methods often struggle with the complexities introduced by camouflaged objects, indicating the need for substantial improvements. By releasing COTD, we aim to advance tracking technology to address the unique challenges posed by camouflaged targets and ultimately enhance the accuracy and reliability of visual tracking systems across various applications. The main contributions of this paper are as follows:
  • We introduce COTD, the first benchmark for tracking camouflaged objects, with 200 meticulously annotated sequences. COTD presents unique challenges not addressed by existing benchmarks and may inspire further research in this area.
  • We thoroughly evaluate 20 state-of-the-art trackers on COTD, highlighting their limitations, setting performance benchmarks, and inspiring future research in tracking camouflaged objects.
  • We introduce HIPTrack-MLS, a novel tracker that leverages multi-level features to enhance performance, establishing a stronger baseline for future research in tracking camouflaged objects.
II Related Works
II-A Visual Tracking Algorithms
DCF-based trackers treat the visual tracking problem as an online regression task, leveraging the Parseval theorem and Fast Fourier Transform (FFT) for efficient computation. The MOSSE (Minimum Output Sum of Squared Error) filter [39] is an early representative in this domain. Since its introduction, significant advancements have been made. For example, [40, 41] introduced additional scale filters to address variations in object scale, while [42] and [43] employed regularization techniques to enhance tracker robustness. Furthermore, extensions of the DCF framework, such as translation-invariant techniques, continue to evolve [44]. However, due to the limited representational capacity of handcrafted features, DCF-based trackers still face challenges in robustness under complex conditions.
To overcome the limitations of DCF-based trackers, deep learning-based trackers have garnered significant attention in recent years. These trackers utilize neural networks to automatically learn features, substantially improving tracking performance. SiamFC [45] pioneered this field by applying Siamese networks [46] to visual tracking. SiamFC learns a nonlinear similarity function through offline training and performs correlation operations on candidate regions and the initial object state to generate response maps for the search area. Building on this, the SiamRPN algorithm introduced a candidate region generation network structure with classification and regression branches to enhance tracking speed. Subsequent research, including SiamRPN++ [47], SiamR-CNN [48], SiamCAR [49], and Ocean [50], has enhanced the accuracy and robustness of Siamese trackers through techniques like attention mechanisms in SiamCAR, backbone optimization in SiamRPN++, ResNet-50 feature enhancement in Ocean, and two-stage detection in SiamR-CNN.
Recently, the emergence of Vision Transformers (ViTs) [51] has provided new directions for the development of visual tracking technologies. Vision Transformers demonstrate potential in simplifying and unifying tracking frameworks [51, 52, 53, 54]. For instance, OStrack [54] integrates feature extraction and relational modeling, and introduces an early elimination module to achieve an efficient single-stream tracking framework. DropTrack [53] improves temporal correspondence learning by adaptively performing spatial attention dropout during frame reconstruction. HIPTrack [55] introduces a history prompting network that utilizes refined historical foreground masks and object visual features, providing comprehensive and precise guidance for trackers.
II-B Visual Tracking Benchmarks
Over the past decade, numerous benchmarks have been proposed for testing and evaluating single-object tracking performance. The advancement of visual tracking technology relies on accurate measurement and evaluation of tracking methods, with benchmark datasets serving as a critical evaluation standard. Currently, there are two main types of benchmark datasets: general benchmarks and specific benchmarks [56].
General benchmarks: General benchmark datasets are primarily used to assess the performance of tracking algorithms in typical scenarios. OTB [57, 58] and VOT [59, 60, 61, 62, 63, 64] are among the initial platforms for unified testing data and evaluation standards for general visual tracking. NFS [65] focuses on high-frame-rate videos. Many large-scale benchmarks have been introduced for training deep learning-based trackers. For example, TrackingNet [66] aims to provide resources for training and evaluating general object tracking algorithms, containing over 30,000 sequences. GOT-10K [67] includes over 10,000 videos from the WordNet semantic hierarchy to develop the generalization capability of trackers. LaSOT [68] encompasses 1,400 long-term video sequences, with an additional 150 sequences and a new evaluation scheme.
Specific benchmarks: Specific benchmark datasets aim to evaluate tracking performance for particular types of objects. For example, UAV123 [69] focuses on low-altitude drone object tracking, containing 123 sequences captured by drones. CDTB [70] and PTB [71] are designed to evaluate tracking performance in RGB-D videos, where D represents depth images. TOTB [56] specializes in tracking transparent objects, collecting 225 videos from 15 categories of transparent objects. Recently, TSFMO [72] has been introduced for tracking small and fast-moving objects, while TRO [73] focuses on tracking reflected objects.Additionally, BihOT [74] contains 49 videos and 42K frames, focusing on hyperspectral camouflage. UW-COT220 [75] focuses on underwater camouflage
Despite the availability of existing benchmarks, they mainly focus on objects that do not intentionally blend into their surroundings. Camouflaged objects, however, present a unique challenge that these benchmarks do not adequately address. To fill this critical gap, we propose the Camouflaged Object Tracking Dataset (COTD) in this work.
II-C Dealing with camouflaged Objects in Vision
Camouflaged objects are difficult to distinguish from their backgrounds because they blend into their surroundings. Unlike regular objects, they have similar textures, colors, and patterns as their environment, creating visual ambiguity. Recently, many computer vision techniques have been developed to tackle these challenges. For instance, Fan et al. [32] introduced the task of Camouflaged Object Detection (COD) and established a dataset COD10K with 10,000 images, along with a simple yet effective framework SINet for efficient and precise detection of camouflaged objects. Liu et al. [33] proposed a scale-adaptive correlation filter algorithm that integrates FHOG and LBP features for Camouflaged Object Segmentation (COS). Yue et al. [76] and Chen et al. [77] have further refined segmentation accuracy by integrating multi-scale feature fusion techniques. Cheng et al. [26] developed a Video Camouflaged Object Detection (VCOD) framework that captures inter-frame motion and uses a spatio-temporal transformer for consistency, along with the MoCA-Mask dataset providing detailed pixel-level annotations. Additional research by Wang et al. [78] have demonstrated the potential of leveraging discriminative feature learning for enhancing detection robustness in complex environments. Despite these advancements, Camouflaged Object Tracking (COT), which focuses on consistently identifying camouflaged objects in videos while adapting to changes in position, scale, and appearance, remains under-explored. The absence of publicly available datasets for camouflaged object tracking has further constrained the progress of research and algorithm development in this field. To bridge this gap, we introduce the Camouflaged Object Tracking Dataset (COTD), specifically designed to support the development and evaluation of algorithms tailored for tracking camouflaged objects.
III Construction of the COTD Dataset
III-A Data Collection
Figure 3: This chart depicts the distribution of sequence counts for 20 categories. Each category is represented by a bubble, where the size of the bubble corresponds to the number of sequences in that category. Larger bubbles indicate categories with higher sequence counts, visually highlighting the differences in quantity among the categories.
To thoroughly assess the performance of camouflaged object tracking algorithms, we develop a novel benchmark dataset known as the Camouflaged Object Tracking Dataset (COTD). COTD offers a diverse and high-quality platform for advancing and evaluating camouflaged object tracking technologies. Our rigorous and extensive data collection process involve extensive internet searches and scientific literature mining, initially gathering hundreds of potential high-quality videos. These videos encompass a variety of environments, including forests, grasslands, oceans, deserts, and snowfields, each carefully selected to exhibit distinct behaviors and motion patterns of camouflaged targets.
In the video selection process, we apply stringent criteria to ensure the high quality and practicality of the data. Specifically, videos are evaluated and selected based on three main criteria: 1) the diversity and complexity of the environment, 2) the frequency with which targets cross frame boundaries, and 3) the overall length and continuity of the videos. We prioritize videos that present additional challenges, such as a high degree of similarity between targets and backgrounds. Due to resource constraints in collecting, processing, and annotating video data, we select 200 sequences for the benchmark test. This ensures that each video is of high quality with comprehensive annotations, adhering to our available time, budget, and manpower constraints. The COTD dataset comprises 20 categories, namely: bird, bug, butterfly, cat, crab, dog, fish, frog, grasshopper, leopard, lizard, mantis, octopus, person, phasmatodea, seahorse, rabbit, shrimp, snake, and spider. To visually represent the composition of the dataset, we create a chart illustrating the number of sequences per category. Additionally, we generate a series of graphs depicting the maximum, minimum, and average number of frames for each category. These visual aids not only enhance the understanding of the dataset’s diversity but also highlight the challenges in tracking camouflaged objects across different scenarios. Fig. 3 presents the distribution of sequences across the 20 categories, while Fig. 4 provides a detailed view of the frame characteristics for each category, such as the range and average number of frames. These figures are crucial in demonstrating the dataset’s comprehensiveness and its potential as a robust benchmark for evaluating the performance of camouflaged object tracking algorithms. Additionally, to highlight the challenges associated with tracking camouflaged objects, we evaluate seven state-of-the-art trackers on the COTD dataset and compare them with two popular generic object tracking benchmarks. The AUC results are presented in Table I. Notably, the AUC of these trackers on the TrackingNet benchmark is more than 15.0% higher than on the COTD dataset, underscoring the significant challenges posed by tracking camouflaged objects.
Compared to existing benchmark datasets, the COTD (ours) dataset focuses on tracking camouflaged objects, which presents a unique challenge in the field of visual single-object tracking. A summarized comparison with existing benchmarks is reported in Table II. The COTD dataset includes 200 videos and 80,000 frames, surpassing in quantity the 100 videos and 59K frames of OTB100 [79], as well as the 70 videos and 15.8K frames of DTB70 [80]. Furthermore, compared to BihOT [74] and UW-COT220 [75], COTD has a higher focus on camouflage tracking. BihOT contains 49 videos and 42K frames, specializing in hyperspectral camouflage, while UW-COT220, although it includes 220 videos and 159K frames, focuses on underwater camouflage. However, the scale and diversity of the COTD dataset enable it to better simulate the complexity and challenges of tracking camouflaged objects in real-world scenarios. COTD specializes in tracking camouflaged objects in various environments, contrasting with the application scenarios of BihOT and UW-COT220.
Figure 4: The radar chart compares video frame counts (minimum, average, maximum) across 20 categories. Blue, orange, and green lines represent the distributions of minimum, average, and maximum frame counts respectively for each category, illustrating differences between them.
TABLE I: AUC (%) Comparison of state-of-the-art trackers on the COTD and generic object tracking benchmarks.
Dataset COTD TrackingNet [66] LaSOT [68]
HIPTrack [81] 68.0 84.5 72.7
ROMTrack [82] 66.3 84.1 71.4
DropTrack [83] 67.4 84.1 71.8
ARTrack [84] 66.7 85.1 72.6
SeqTrack [85] 64.1 83.9 71.5
GRM [86] 64.8 84.0 69.9
SimTrack [87] 63.9 83.4 70.5
TABLE II: Comparison of current popular benchmarks for visual single object tracking in the literature. $\ast$ indicates that only the test set of each dataset is reported.
Benchmark Year Videos Total frames classes Attr. Avg.frames Resolution Visual annotation Feature
OTB100 [79] 2015 100 59K 16 11 590 128$\times$96 $\sim$ 800$\times$336 BBox short-term
UAV123 [88] 2016 123 113K 9 12 915 720$\times$480 $\sim$ 1280$\times$720 BBox UAV
DTB70 [80] 2017 70 15.8K - 11 225 1280$\times$720 $\sim$ 1280$\times$720 BBox UAV
UAVDT [89] 2018 50 37.1K 3 8 742 960$\times$540 $\sim$ 1024$\times$540 BBox UAV
UAVTrack112 [90] 2021 112 100K 32 13 895.65 1280$\times$720 $\sim$ 1920$\times$1080 BBox UAV
WebUAV3M [91] 2022 4500 3.3M 223 17 710 608$\times$320 $\sim$ 1982$\times$1088 BBox UAV
TrackingNet* [66] 2018 511 226K 27 15 441.47 270$\times$360 $\sim$ 1280$\times$720 BBox large scale
LaSOT* [68] 2019 280 685K 85 14 2447.71 202$\times$360 $\sim$ 1280$\times$720 BBox category balance
GOT-10k* [67] 2019 180 23K 84 6 133.33 270$\times$480 $\sim$ 3840$\times$2160 BBox generic
VOT [92] 2016 60 21K 37 9 357.58 320$\times$180 $\sim$ 1280$\times$720 BBox annual
TSFMO250 [72] 2022 250 51K 26 12 202.72 720$\times$404 $\sim$ 1310$\times$720 BBox small and fast moving
TRO [73] 2024 200 70K - 12 349.03 244$\times$578 $\sim$ 2400$\times$1080 BBox reflected
LLOL [93] 2024 269 133K 32 12 493.66 544$\times$960 $\sim$ 3840$\times$2160 BBox low-light
MoCA-Mask [26] 2022 87 23K 45 - 263.67 480$\times$360 $\sim$ 1920$\times$1080 Mask camouflaged
BihoT [74] 2024 49 42K - 9 855.35 - BBox hyperspectral camouflaged
UW-COT220 [75] 2025 220 159K 96 - 722 320$\times$240 $\sim$ 3840$\times$2160 BBox+Mask underwater camouflaged
COTD (Ours) 2025 200 80K 20 12 406.55 270$\times$480 $\sim$ 2488$\times$1152 BBox camouflaged
(a) Tight bounding box on the first
frame.
(b) The fish is partially occluded by the hand.
(c) The fish is fully occluded by the hand.
(d) The fish partially passed the hand.
Figure 5: Illustration of our annotation rules using fish-19, where the fish performs a jump and lands on the hand. In (a), the bounding box is kept as tight as possible around the fish before it jumps. In (b) and (d), the fish is partially obscured by the person’s hand. We provide tight bounding boxes that closely follow the visible parts of the fish. In (c), the fish is fully occluded by the hand, and thus no bounding box is provided due to the Full Occlusion (FO) scenario.
III-B Data Annotation
We follow the principles in [66] for sequence annotation. Each frame undergoes meticulous annotation by expert annotators, comprising students specializing in visual tracking research. The annotation process involves drawing axis-aligned bounding boxes that represent the most compact boundary of the target object, with careful documentation of object visibility, occlusion, and absence states. Examples of bounding box annotations in the COTD are shown in Fig. 2.
To ensure annotation quality and consistency, we implement a rigorous two-stage verification process. In the primary stage, annotators conduct frame-by-frame labeling, ensuring precise object identification even in complex scenes. This is followed by a secondary verification where another annotator reviews all annotations, with any discrepancies resolved through discussion and consensus.
The details of our annotation rules are as follows.
Rule 1: All visible components of the designated object must be encompassed within the bounding box, which should be drawn as tightly as possible around the object.
Rule 2: In instances where the target object is entirely occluded or exits the frame, no bounding box should be provided for that particular frame.
Rule 3: A valid annotated bounding box must be present in the initial frame of each video sequence to delineate the target object.
Rule 4: It is imperative to document the entire gradual transition of the target object when it undergoes occlusion, disappearance, or re-emergence within the scene.
As illustrated in Fig. 5, we adhere to the rules by drawing bounding boxes that are tightly fitted within the boundaries of the target object, avoiding overly large sizes to ensure precision across numerous frames. Additionally, we focus on accurately annotating the processes of the target object’s disappearance and re-emergence (either FO or OV), which is crucial for training trackers to handle occlusions and the recapture of targets. This meticulous annotation helps create specific test scenarios, such as long-term occlusions, which are very useful for evaluating the recapture capabilities of tracking algorithms.
III-C Attributes
Figure 6: Distribution of sequences with each attribute in COTD.
Figure 7: The overall framework of HIPTrack-MLS, which is inherited from HIPTrack. The primary difference is in the encoder component of the feature extraction network. The original one is replaced by the proposed Multi-Level Feature Sensing Encoder.
Figure 8: Structure of the Multi-Level Feature Sensing Encoder.
Following the prevalent tracking benchmark methodologies [94], [67], and [58], we meticulously annotate each video sequence with detailed attributes to enable an in-depth analysis of tracking algorithms’ performance. The proposed COTD dataset features 12 attributes, with each sequence including Background Clutter (BC) to highlight the dataset’s camouflage characteristics. In addition, we annotate the following 11 attributes: (1) Illumination Variation (IV), (2) Scale Variation (SV), (3) Deformation (DEF), (4) Motion Blur (MB), (5) Fast Motion (FM), which is marked when the object center moves by at least 20% of its own size between consecutive frames, (6) Out-of-View (OV), (7) Low Resolution (LR), marked when the object area is smaller than 900 pixels, (8) Partial Occlusion (POC), (9) Rotation (ROT), (10) Full Occlusion (FOC), and (11) Aspect Ratio Change (ARC), marked when the object’s bounding box aspect ratio falls outside the range [0.5, 2]. Fig.6 shows the distribution of these attributes on COTD. As can be seen, the most common challenge in COTD is Motion Blur. In addition, the Deformation and Partial Occlusion also present frequently in COTD.
TABLE III: Co-occurrence distributions of the attributes in COTD. The diagonal demonstrates the distribution of the entire dataset, and each row or column represents the distribution of the corresponding attribute subset.
Attribute IV SV DEF MB FM OV LR POC ROT FOC ARC
IV 5 0 1 1 0 0 0 2 0 0 0
SV 0 29 14 12 7 8 3 17 2 1 4
DEF 1 14 59 32 9 16 5 27 13 3 8
MB 1 12 32 72 12 19 10 25 8 4 6
FM 0 7 9 12 14 5 2 8 3 2 1
OV 0 8 16 19 5 39 3 12 4 2 2
LR 0 3 5 10 2 3 21 6 0 0 0
POC 2 17 27 25 8 12 6 58 11 4 6
ROT 0 2 13 8 3 4 0 11 29 1 3
FOC 0 1 3 4 2 2 0 4 1 4 0
ARC 0 4 8 6 1 2 0 6 3 0 9
Based on the 11 attributes, the co-occurrence matrix of the COTD dataset is provided in Table III. The horizontal and vertical axes correspond to each attribute, and the numbers in the table refer to the number of sequences with both horizontal and vertical attributes. For example, the number of sequences with both MB and DEF is 32. By comprehensively considering these attributes, we can conduct a thorough evaluation of tracking algorithms under various challenging conditions.
IV Method
In our evaluation, we observed that while HIPTrack [81] outperforms other advanced trackers in COTD, its performance still falls significantly short of satisfactory levels. Camouflaged objects often share similar textures, colors, and patterns with their environment, creating a high level of visual ambiguity. This makes them challenging to detect using single-level feature representation. To overcome this, we propose a novel baseline tracker based on HIPTrack, called HIPTrack-MLS, which enable HIPTrack to sense multi-level features to tackle the difficulties posed by camouflaged objects. By capturing information at various scales and resolutions, multi-level features enhance HIPTrack’s ability to distinguish targets from their environment. This adaptation significantly improves the tracker’s representation power, increasing the chances of accurately identifying camouflaged objects.
Figure 9: Overall performance on COTD. Precision and success rate for one-pass evaluation (OPE) [57] are used for evaluation.
IV-A Overall Architecture
Similar to HIPTrack [81], HIPTrack-MLS comprises three main components, as illustrated in Fig. 7 and labeled as (a), (b), and (c), respectively. The feature extraction network is responsible for extracting features from the search region containing target-template matching information while filtering out background image patches. The historical prompt network includes a historical prompt encoder and a historical prompt decoder. The historical prompt encoder encodes the position information and visual features of the target in the current frame as historical target features and stores them in the memory bank of the historical prompt decoder. The historical prompt decoder generates historical prompts and combines them with the compressed features of the search region. The head network has the same structure as OSTrack [95]. The key difference between HIPTrack-MLS and HIPTrack lies in their feature extraction networks. While HIPTrack uses a Vision Transformer (ViT) with early candidate elimination modules, similar to the feature encoder used in OSTrack, HIPTrack-MLS is characterized by a novel Multi-Level Feature Sensing Encoder architecture, as illustrated in Fig. 8. It is designed to enhance feature representation capabilities for camouflaged objects by perceiving features across multiple levels of abstraction. The following section will provide a detailed description of this proposed encoder. Details of other architectures inherited from HIPTrack can be found in [81].
IV-B Multi-Level Feature Sensing Encoder
Our Multi-Level Feature Sensing Encoder is an extension of the Transformer Encoder architecture used in HIPTrack, specifically a Vision Transformer (ViT) equipped with early candidate elimination modules. This structure is comprised of stacked CEBlocks, as illustrated in Fig. 8. The main distinction is the introduction of a Multi-Level Feature Sensing structure for processing multi-level features output by the ViT blocks. In this updated design, features from all levels are directly combined through concatenation, allowing the model to use both detailed information from lower levels and abstract insights from higher levels. A multilayer perceptron efficiently reduces the dimensionality of these multi-level features, using three linear layers with ReLU activation functions to compress and refine the feature representation. Through a process of reshaping and permutation, the module adapts the data to the linear layers, sequentially reducing the dimensionality from 12 to 6, then to 3, and finally to 1. This series of reductions is supplemented by ReLU activation, introducing non-linearity and enhancing the model’s ability to capture complex feature patterns. Given that the early candidate elimination module in each CEBlock might alter the input feature’s shape, we split the CEBlock into two parts: one representing the ViT block with early candidate elimination ($\mathbf{CBlock}^{\rightarrow}_{i}$) and the other without it ($\mathbf{CBlock}^{\downarrow}_{i}$). The corresponding output features are denoted by $H^{\rightarrow}_{i}$ and $H^{\downarrow}_{i}$, respectively. The effect of our proposed Multi-Level Feature Sensing Encoder can be expressed as:
${F}_{mls}=(1-\gamma)\mathfrak{P}\left(H_{n}^{\rightarrow}\right)+\gamma\text{MLP}\left([H^{\downarrow}_{1},...,H^{\downarrow}_{n}]\right),$ (1)
where $H^{\downarrow}_{i}=\mathbf{CBlock}^{\downarrow}_{i}(H^{\downarrow}_{i-1})$, $H^{\rightarrow}_{i}=\mathbf{CBlock}^{\rightarrow}_{i}(H^{\rightarrow}_{i-1})$, $\mathfrak{P}$ represents padding and reshaping used to align $H^{\rightarrow}_{n}$ with $H_{0}$, and $\gamma\in[0,1]$ is a weighting constant balancing the importance of the aggregated feature and $H^{\rightarrow}_{n}$. $\mathbf{MLP}$ denotes the multilayer perceptron network. When $\gamma=0$, $F_{mls}$ is simply the feature produced by the Transformer Encoder of HIPTrack. Thus, our Multi-Level Feature Sensing Encoder extends HIPTrack’s feature extraction network, enhancing feature representation by capturing information at multiple levels of abstraction. Since the proposed encoder does not introduce additional training loss, we can use the same training pipeline as HIPTrack for training HIPTrack-MLS. For details, please refer to HIPTrack [81].
V Evaluation
Figure 10: Attribute-based comparison on Aspect Ration Change, Illumination Variation, Out-of-View, and Partial Occlusion.
In this study, precision and success rate serve as the quantitative measures for evaluation.
Precision plot: The precision plot is a standard metric used for evaluating tracking performance. It assesses tracking accuracy by measuring the center position error, which calculates the average Euclidean distance between the tracked object’s center and its ground truth position. This metric summarizes the tracking performance by averaging the errors across all frames in a sequence. However, when a tracker loses the object, its output may become arbitrary, reducing the reliability of the average error [96]. To mitigate this issue, the precision plot is commonly used. It evaluates the overall tracking accuracy by comparing the tracker’s estimated position to the ground truth within a specified distance threshold [96, 97]. In our evaluation, we use a 20-pixel threshold for the precision score, in line with established practices [96].
Success plot: Another evaluation metric commonly used is the success plot. This metric gauges tracker performance by analyzing the overlap between predicted and ground truth bounding boxes. Specifically, for a given tracking bounding box $r_{t}$ and a ground truth bounding box $r_{a}$ , the overlap score is computed as $S=\frac{\left|r_{t}\bigcap r_{a}\right|}{\left|r_{t}\bigcup r_{a}\right|}$ , where $\cap$ and $\cup$ denote the intersection and union of the two bounding boxes, respectively, and ||$\cdot$|| ⋅ | represents the number of pixels contained within these regions. To assess tracking performance across a sequence of frames, we count the number of frames where the overlap score $S$ exceeds a predefined threshold $t_{o}$. The success plot then illustrates the proportion of frames deemed successful at various overlap thresholds ranging from 0 to 1. This visual representation helps in understanding how often the tracker achieves sufficient overlap with the ground truth over different levels of thresholding.Relying on a single success rate at a specific threshold (e.g. $t_{o}$=0.5) may not offer a comprehensive view of a tracker’s performance, as it provides a limited perspective on tracking accuracy. To address this, the area under the curve (AUC) is typically employed in conjunction with the success plot to rank tracking algorithms [96, 97]. The AUC aggregates the success rate across all threshold values, offering a thorough evaluation of the tracking algorithm’s performance across various conditions. This method facilitates a more nuanced and accurate comparison of different tracking algorithms by providing a holistic view of their effectiveness in various scenarios. Each tracker is evaluated using the default parameters provided in the source code. We employ one-pass evaluation (OPE) and assess precision and success rate, as per [98, 99].
V-A Trackers for Comparison
We evaluate our HIPTrack-MLS and 20 state-of-the-art trackers on COTD, including HIPTrack [81], ROMTrack [82], SeqTrack [85], DropTrack [83], AQATrack [100], SparseTT [101], DCPT [102], GRM [86], Stark [103], ZoomTrack [104], MAT [105], SimTrack [87], ARTrack [84], CSWinTT [106], JointNLT [107], TCTrack [108], SiamGAT [109], UDAT [110], HiFT [111], and Aba-ViTrack [112].
V-B Evaluation Results
Overall performance: A comprehensive evaluation was conducted on 20 state-of-the-art trackers and our HIPTrack-MLS using the COTD dataset. The evaluation results are shown in Fig. 9. Notably, HIPTrack-MLS achieves superior performance compared to all other trackers, with a precision (PRC) of 0.525, which is 1.0% higher than its baseline, HIPTrack, which scored 0.515. In terms of success rate (AUC), HIPTrack-MLS also leads with an AUC of 0.692, surpassing the second-place HIPTrack, which has an AUC of 0.680, by 1.2%. These results highlight the outstanding performance of HIPTrack-MLS in both precision and success rate metrics, showcasing its superiority in tackling the challenges of tracking camouflaged objects. This success is largely due to HIPTrack-MLS’s advanced multi-level feature Sensing, which effectively leverages multi-level features.
Figure 11: Qualitative evaluation on 4 sequences from COTD, i.e. fish-4, mantis-2, person-9, and phasmatodea-10 from top to bottom. The results of different methods have been shown with different colors, and ‘GT’ denotes the ground truth.
Attribute-based performance: We conducte a comprehensive performance evaluation on eleven common attributes to better analyze and understand the performance of different trackers on the COTD dataset. HIPTrack-MLS consistently ranks first across most attribute subsets. Due to paper length constraints, we present the PRC results for four common challenges in Fig. 10, including Aspect Ration Change (ARC), Illumination Variation (IV), Out-of-View (OV), and Partial Occlusion (POC). As shown, HIPTrack-MLS consistently ranks first across all these attribute subsets, surpassing the baseline HIPTrack by at least 1.0% in all these subsets. Notably, HIPTrack-MLS achieves a precision of 0.271 in the IV subset, surpassing the second place HIPTrack by 3.4%. For the ARC and OV attributes, HIPTrack-MLS shows improvements of 1.3% and 1.0%, respectively, over HIPTrack.
Evaluation on different object categories: In our research, we meticulously conduct a comprehensive experiment to evaluate the performance of multiple state-of-the-art trackers when handling camouflaged objects, with the results systematically presented in Table IV. We strategically categorize the COTD dataset into two superclasses—Aquatic and Non-aquatic—to thoroughly test the capabilities of 20 state-of-the-art trackers, alongside our tracker, HIPTrack-MLS. The experiment hinges on three pivotal metrics: AUC, Precision, and Normalized Precision, to provide a robust assessment of each tracker’s performance. The experimental results demonstrate that our improved tracker HIPTrack-MLS excels in all evaluated metrics, significantly outperforms its baseline, HIPTrack. Specifically, in the domain of non-aquatic object tracking, HIPTrack-MLS achieves an AUC that is 0.58% higher than HIPTrack, with an increase from 63.95% to 64.53%. Additionally, HIPTrack-MLS also leads with a precision of 44.52%, surpassing the second-place HIPTrack, which has a precision of 43.69%, by 0.83%. While there is a minor 0.12% reduction in normalized precision, from 74.22% to 74.10%, this decrement is negligible in the context of real-world applications. When it comes to the aquatic objects, HIPTrack-MLS with a AUC of 77.21%, which is 1.79% higher than its baseline, HIPTrack, which scores 75.42%. For the precision and normalized precision, HIPTrack-MLS shows improvements of 1.08% and 1.48%, respectively, over HIPTrack. These results indicate that our tracker performs exceptionally well in the tracking of camouflaged objects. Additionally, we observe performance variances among other trackers in both objects. Some trackers, such as AQATrack and HIPTrack, show higher precision and normalized precision in aquatic objects, while others, like HiFT and SiamGAT, exhibit relatively lower AUC in non-aquatic objects. This observation potentially indicates limitations in their generalization capabilities and underscores the importance of further research in this area.
TABLE IV: Overall tracking results of the trackers on Non-aquatic and Aquatic,The first-, second- and third-place trackers are labeled with red, blue and green colors respectively.
Tracker Venue Non-aquatic Aquatic
AUC Pre nPre AUC Pre nPre
HiFT [111] ICCV 2021 49.95 29.38 56.19 54.84 25.69 66.90
STARK [103] ICCV 2021 57.93 35.01 66.42 69.80 42.16 84.58
SiamGAT [109] CVPR 2021 50.73 24.45 59.11 59.02 21.79 74.60
SimTrack [87] ECCV 2022 60.02 36.44 67.82 71.01 46.22 85.36
SparseTT [101] IJCAI 2022 57.87 36.00 68.17 69.13 40.93 82.66
CSWinTT [106] CVPR 2022 56.67 32.04 64.04 68.56 43.10 81.76
TCTrack [108] CVPR 2022 38.52 22.68 45.85 46.37 20.87 59.28
UDAT [110] CVPR 2022 49.15 29.61 58.10 59.73 32.75 73.93
Aba-ViTrack [112] ICCV 2023 56.34 33.48 64.80 67.91 40.04 81.56
ROMTrack [82] ICCV 2023 61.16 40.59 71.77 74.90 58.47 87.46
ARTrack [84] CVPR 2023 61.18 40.43 69.80 75.76 58.42 88.83
DropTrack [83] CVPR 2023 63.03 42.61 73.08 75.03 56.14 87.65
GRM [86] CVPR 2023 59.59 38.29 69.06 73.29 53.99 86.56
JointNLT [107] CVPR 2023 44.41 19.94 50.46 58.99 27.45 70.95
MAT [105] CVPR 2023 56.86 33.47 64.35 64.92 37.25 78.03
SeqTrack [85] CVPR 2023 59.23 40.63 70.22 72.29 53.03 86.24
ZoomTrack [104] NeurIPS 2023 58.72 36.86 67.88 72.58 53.72 85.48
DCPT [102] ICRA 2024 60.27 38.71 68.32 73.41 52.04 86.75
AQATrack [100] CVPR 2024 61.74 41.73 71.53 75.72 56.22 88.82
HIPTrack [81] CVPR 2024 63.95 43.69 74.22 75.42 56.42 88.01
HIPTrack-MLS Ours 64.53 44.52 74.10 77.21 57.50 89.49
TABLE V: Efficiency comparison evaluated on COTD.
HIPTrack-MLS HIPTrack ProContEXT ZoomTrack DCPT
FPS 15.67 23.46 10.87 15.62 11.25
TABLE VI: Performance comparison of HIPTrack-MLS and its baseline HIPTrack on generic and our benchmarks.
Tracker LaSOT [98] TrackingNet [99] COTD
Prec. Succ. Prec. Succ. Prec. Succ.
HIPTrack [55] 0.795 0.727 0.838 0.845 0.515 0.680
HIPTrack-MLS 0.793 0.726 0.837 0.845 0.525 0.692
Efficiency analysis: To evaluate the computational efficiency of our proposed method, we conducted runtime analysis on the COTD benchmark dataset. Table Table V presents the tracking speed in frames per second (FPS) for HIPTrack-MLS and other state-of-the-art trackers. The results show that HIPTrack-MLS achieves a tracking speed of 15.67 FPS, which is comparable to ZoomTrack (15.62 FPS) and faster than ProContEXT (10.87 FPS) and DCPT (11.25 FPS). While the original HIPTrack demonstrates higher efficiency at 23.46 FPS, our multi-level feature sensing architecture introduces additional computational overhead due to the feature fusion and dynamic weighting mechanisms. However, this trade-off between speed and accuracy is justified by the significant performance improvements in tracking camouflaged objects, as demonstrated in our previous experiments.
Performance on generic datasets: To validate whether HIPTrack-MLS’s improvements are specifically beneficial for camouflaged object tracking or represent a general enhancement in tracking capability, we conduct comprehensive evaluations on standard tracking benchmarks including LaSOT [98] and TrackingNet [99], alongside our COTD dataset. The results are presented in Table VI. On the LaSOT dataset, HIPTrack-MLS achieves comparable performance to the baseline HIPTrack, with precision and success scores of 0.793 and 0.726, respectively, showing a slight marginal difference compared to HIPTrack’s 0.795 and 0.727. Similarly, on TrackingNet, HIPTrack-MLS demonstrates minimal improvement with precision and success scores of 0.837 and 0.845, compared to HIPTrack’s 0.838 and 0.845. However, on the COTD dataset, HIPTrack-MLS shows a significant performance gain, improving the success rate from 0.680 to 0.692 and precision from 0.515 to 0.525. This notable improvement on COTD, coupled with maintained performance on general tracking benchmarks, suggests that our multi-level feature sensing architecture specifically addresses the challenges of camouflaged object tracking without compromising general tracking capability.
The experimental results support our hypothesis that the multi-level feature sensing mechanism is particularly effective for camouflaged objects, where fine-grained feature discrimination at different scales is crucial for maintaining tracking accuracy. The maintained performance on general tracking benchmarks also indicates that our modifications do not introduce any degradation in general tracking scenarios, making HIPTrack-MLS a versatile tracking solution.
Qualitative evaluation: In Fig. 11, we demonstrate qualitative comparisons of our method with 9 top trackers, including HIPTrack [81], DropTrack [83], DCPT [102], AQATrack [100], ARTrack [84], GRM [86], ROMTrack [82], SeqTrack [85], and ZoomTrack [104]. These examples face challenges such as Illumination Variation, Deformation, Rotation, Aspect Ratio Change, and Partial Occlusion. Specifically, in the fish-4 sequence, only HIPTrack-MLS consistently tracks the camouflaged fish. In mantis-2, while some trackers manage to track the mantis, HIPTrack-MLS consistently demonstrates superior visual performance. In person-9, due to background clutter and multiple object interference, only HIPTrack-MLS, DropTrack, ARTrack, and SeqTrack can track the target person. In phasmatodea-10, many trackers are affected by the background, leading to inaccurate tracking of the camouflaged phasmatodea. These qualitative results validate the superiority of our method for tracking camouflaged objects.
TABLE VII: Illustrates the impact of the importance coefficient of aggregated feature on HIPTrack-MLS’s PRC (%) and AUC (%) on COTD. Red, blue, and green represent the first, second, and third place, respectively.
$\gamma$ 1.0 0.9 0.8 0.7 0.6 0.5 0.4 0.3 0.2 0.1
PRC (%) 47.0 48.1 49.1 49.9 50.4 51.1 51.7 52.1 52.2 52.5
AUC (%) 66.7 67.4 68.0 69.1 68.1 68.2 68.7 68.8 68.3 69.2
Figure 12: For each group, the top shows the original images, while the following two rows display the attention map generated by the algorithm in some video frames. The middle row is from HIPTrack, and the bottom row is from HIPTrack-MLS.
V-C Ablation study
Impact of the importance coefficient in Multi-Level Feature Sensing. To examine the effect of the weight coefficient $\gamma$ on balancing the contributions of the aggregated multi-level features and $H_{n}^{\rightarrow}$ in HIPTrack-MLS, we evaluated HIPTrack-MLS on the COTD dataset with a range of $\gamma$ settings. Specifically, $\gamma$ ranges from 0.1 to 1.0 with increments of 0.1. As $\gamma$ increases, the influence of the aggregated feature becomes more pronounced, while the reliance on $H_{n}^{\rightarrow}$ decreases. Table VII details HIPTrack-MLS’s PRC and AUC on the COTD dataset for various $\gamma$ values. The results indicate that the optimal PRC and AUC, which are 52.5% and 69.2% respectively, are achieved when $\gamma$ is set to 0.1. This result indicates that an appropriate choice of $\gamma$ enhances tracking performance, confirming the effectiveness of the proposed method. Consequently, we adopted $\gamma=0.1$ as the default setting for HIPTrack-MLS.
Attention mechanism analysis. To make sense of the advantage of the propose method, we explore the allocation of attention in the task of Camouflaged Object Tracking and compare the performance of the improved HIPTrack-MLS model with the original HIPTrack model. Fig. 12 intuitively displays three types of visual representations: the original image, the attention map of the original HIPTrack model, and the attention map of the HIPTrack-MLS model. The original images in the first row demonstrate the camouflage of the object in a complex background, which poses a significant challenge to tracking algorithms. The high degree of visual integration between the object and the background increases the difficulty of tracking. The second row shows the attention maps generated by the original HIPTrack model. These maps reveal some limitations of the original model when handling camouflaged objects, especially when the similarity between the object and the background is high, the original model may struggle to maintain precise tracking of the object. The third row, however, showcases the attention maps of the HIPTrack-MLS model, which clearly highlight the target object. Even when the object is camouflaged, the model can accurately identify and focus attention on the key features of the object. This improved attention mechanism significantly enhances tracking accuracy, especially in complex environments.
VI Conclusion
In this work, we delve into the relatively underexplored task of tracking camouflaged objects, a field with significant potential for diverse applications. However, there is currently no publicly available benchmark specifically designed for camouflaged object tracking. To address this gap, we introduce the COTD, which is, to the best of our knowledge, the first benchmark specifically designed for camouflaged object tracking. In addition, to evaluate the performance of existing trackers and establish a benchmark for future research, we conducted a comprehensive evaluation and detailed analysis of 20 state-of-the-art tracking algorithms. Additionally, we introduced a new tracker, HIPTrack-MLS, which enhances HIPTrack by improving its multi-level feature sensing capability for better feature representation. Experimental results demonstrate that HIPTrack-MLS significantly outperforms the current leading algorithms on COTD.